# Sum of Squares Decompositions for Structured Biquadratic Forms

### v1

@@ 5.3 Geometry of the $3\times 3$ PSD Cone
Let $P$ be a $3\times 3$ monic symmetric biquadratic form as in Theorem 5.2. The set of coefficients $(a,b,c)$ for which $P$ is positive semi-definite is a convex polyhedron in $\mathbb{R}^{3}$ described by the four linear inequalities:
$1+2a+2b+4c\geq 0,$ (7)
$1+2a-b-2c\geq 0,$ (8)
$1+2b-a-2c\geq 0,$ (9)
$1-a-b+c\geq 0.$ (10)
The polyhedron has four faces, six edges, and four vertices. Its boundary consists of the planar faces where each inequality becomes an equality.
@@
Given that $c\neq 0$, $\frac{a}{c}\in[-2,1]$, $\frac{b}{c}\in[-2,1]$, and conditions (7)–(10) hold, it follows that $1-\frac{ab}{c}\geq 0$ is always true. Therefore, the psd condition is equivalent to the four linear inequalities (7)–(10). These define a convex polyhedron in $\mathbb{R}^{3}$. The faces are obtained by setting each inequality to equality while the others remain inequalities. The edges and vertices are intersections of these faces.
The explicit description of faces, edges, and vertices can be obtained by solving the linear system. For brevity, we omit the full enumeration here, but note that the polyhedron is bounded (unlike the outer approximation from the three conditions alone). ∎
5.4 Examples on the Boundary of the PSD Cone
As an illustration of Theorem 5.3, we present six explicit monic symmetric biquadratic forms, one on each edge of the psd cone. For each edge, we choose a simple interior point. The faces generated by conditions (7)–(10) are denoted by $F_{1},\dots,F_{4}$, respectively.
  1. 1.
    On edge $E_{12}$ (intersection of $F_{1}$ and $F_{2}$): $a=-\frac{1}{2}$, $b=0$, $c=0$:
    $P_{1}(\mathbf{x},\mathbf{y})=\sum_{i=1}^{3}\sum_{j=1}^{3}x_{i}^{2}y_{j}^{2}-\frac{1}{2}\sum_{i\neq k}\sum_{j=1}^{3}x_{i}x_{k}y_{j}^{2}.$
  2. 2.
    On edge $E_{13}$ (intersection of $F_{1}$ and $F_{3}$): $a=0$, $b=-\frac{1}{2}$, $c=0$:
    $P_{2}(\mathbf{x},\mathbf{y})=\sum_{i=1}^{3}\sum_{j=1}^{3}x_{i}^{2}y_{j}^{2}-\frac{1}{2}\sum_{i=1}^{3}\sum_{j\neq l}x_{i}^{2}y_{j}y_{l}.$
  3. 3.
    On edge $E_{14}$ (intersection of $F_{1}$ and $F_{4}$): $a=0$, $b=\frac{1}{2}$, $c=-\frac{1}{2}$:
    $P_{3}(\mathbf{x},\mathbf{y})=\sum_{i=1}^{3}\sum_{j=1}^{3}x_{i}^{2}y_{j}^{2}+\frac{1}{2}\sum_{i=1}^{3}\sum_{j\neq l}x_{i}^{2}y_{j}y_{l}-\frac{1}{2}\sum_{i\neq k}\sum_{j\neq l}x_{i}y_{j}x_{k}y_{l}.$
  4. 4.
    On edge $E_{23}$ (intersection of $F_{2}$ and $F_{3}$): $a=0$, $b=0$, $c=\frac{1}{2}$:
    $P_{4}(\mathbf{x},\mathbf{y})=\sum_{i=1}^{3}\sum_{j=1}^{3}x_{i}^{2}y_{j}^{2}+\frac{1}{2}\sum_{i\neq k}\sum_{j\neq l}x_{i}y_{j}x_{k}y_{l}.$
  5. 5.
    On edge $E_{24}$ (intersection of $F_{2}$ and $F_{4}$): $a=0$, $b=1$, $c=0$:
    $P_{5}(\mathbf{x},\mathbf{y})=(x_{1}^{2}+x_{2}^{2}+x_{3}^{2})(y_{1}+y_{2}+y_{3})^{2}.$
  6. 6.
    On edge $E_{34}$ (intersection of $F_{3}$ and $F_{4}$): $a=1$, $b=0$, $c=0$:
    $P_{6}(\mathbf{x},\mathbf{y})=(x_{1}+x_{2}+x_{3})^{2}(y_{1}^{2}+y_{2}^{2}+y_{3}^{2}).$
These six forms are positive semi-definite and lie on the boundary of the psd cone (since they satisfy two equalities among the defining inequalities).
5.5 SOS Decompositions for Boundary Forms
We have explicit SOS representations for $P_{5}$ and $P_{6}$:
@@ 5.5 SOS Decompositions for Boundary Forms
The forms $P_{5}$ and $P_{6}$ admit explicit representations as sums of squares of bilinear forms:
$P_{5}(\mathbf{x},\mathbf{y})$ $=(x_{1}y_{1}+x_{1}y_{2}+x_{1}y_{3})^{2}+(x_{2}y_{1}+x_{2}y_{2}+x_{2}y_{3})^{2}$
$\quad+(x_{3}y_{1}+x_{3}y_{2}+x_{3}y_{3})^{2},$
$P_{6}(\mathbf{x},\mathbf{y})$ $=(x_{1}y_{1}+x_{2}y_{1}+x_{3}y_{1})^{2}+(x_{1}y_{2}+x_{2}y_{2}+x_{3}y_{2})^{2}$
$\quad+(x_{1}y_{3}+x_{2}y_{3}+x_{3}y_{3})^{2}.$
Thus, $P_{5}$ and $P_{6}$ are SOS with sos rank at most three.
Numerical experiments via semidefinite programming confirm that $P_{1}$, $P_{2}$, $P_{3}$, and $P_{4}$ are also SOS. In particular, the form $P_{4}$ has been verified to be SOS, and its matrix representation $M$ has rank 5, suggesting an sos rank of at most 5.
Consider also the following example, which lies in the interior of the PSD cone:
$P_{0}(\mathbf{x},\mathbf{y})=(x_{1}^{2}+x_{2}^{2}+x_{3}^{2})(y_{1}^{2}+y_{2}^{2}+y_{3}^{2}).$
This $3\times 3$ biquadratic form is a monic symmetric biquadratic form with $a=b=c=0$. It is psd and sos, and can be expressed as the sum of four squares of bilinear forms:
$P_{0}(\mathbf{x},\mathbf{y})=(x_{1}y_{1}+x_{2}y_{2}+x_{3}y_{3})^{2}+(x_{1}y_{2}-x_{2}y_{1})^{2}+(x_{1}y_{3}-x_{3}y_{1})^{2}+(x_{2}y_{3}-x_{3}y_{2})^{2}.$
This decomposition is minimal in the sense that the sos rank of $P_{0}$ is exactly four. Thus, any upper bound on the sos rank for the class of symmetric psd biquadratic forms must be at least four.
@@
Every $3\times 3$ psd monic symmetric biquadratic form $P$ defined by (5) is a sum of squares of bilinear forms. Consequently, every $3\times 3$ psd symmetric biquadratic form is a sum of squares of bilinear forms.
@@
By Theorem 5.3, the set of coefficients $(a,b,c)$ for which the monic symmetric biquadratic form $P$ is PSD is a bounded convex polyhedron (a polytope) with vertices $V_{1}=\left(-\tfrac{1}{2},\ -\tfrac{1}{2},\ \tfrac{1}{4}\right)$, $V_{2}=\left(-\tfrac{1}{2},\ 1,\ -\tfrac{1}{2}\right)$, $V_{3}=\left(1,\ -\tfrac{1}{2},\ -\tfrac{1}{2}\right)$, and $V_{4}=\left(1,\ 1,\ 1\right).$ We first show that the forms corresponding to these vertices are SOS.
For $V_{4}$, we have $P_{V_{4}}(\mathbf{x},\mathbf{y})=\left(\sum_{i=1}^{3}x_{i}\right)^{2}\left(\sum_{j=1}^{3}y_{j}\right)^{2}$, which is the square of the bilinear form $(\sum_{i}x_{i})(\sum_{j}y_{j})$, hence SOS.
For $V_{1}$, we have the identity
$P_{V_{1}}(\mathbf{x},\mathbf{y})=\frac{1}{4}\left(\sum_{1\leq i<j\leq 3}(x_{i}-x_{j})^{2}\right)\left(\sum_{1\leq j<l\leq 3}(y_{j}-y_{l})^{2}\right),$
which can be expanded as
$P_{V_{1}}(\mathbf{x},\mathbf{y})=\frac{1}{4}\sum_{1\leq i<j\leq 3}\sum_{1\leq k<l\leq 3}\bigl((x_{i}-x_{j})(y_{k}-y_{l})\bigr)^{2},$
is a sum of squares of bilinear forms.
For $V_{2}$ and $V_{3}$, by symmetry it suffices to consider $V_{2}$. For $V_{2}$, we have
$P_{V_{2}}(\mathbf{x},\mathbf{y})=\frac{1}{2}\sum_{i=1}^{m}\left(x_{i}\sum_{j=1}^{n}y_{j}\right)^{2}+\frac{1}{2}\sum_{1\leq i<k\leq m}\left((x_{i}-x_{k})\sum_{j=1}^{n}y_{j}\right)^{2}$
is a sum of squares of bilinear forms. The same argument applies to $V_{3}$ by symmetry.
Thus all vertices of the PSD polytope are SOS. Since the polytope is the convex hull of its vertices and the set of SOS forms is a convex cone, every point in the polytope corresponds to an SOS form. This proves the first statement.
Now let $Q$ be an arbitrary $3\times 3$ symmetric biquadratic form that is PSD. By the symmetry, all its diagonal coefficients $q_{ijij}$ are equal. If they are all zero, then $Q$ is identically zero (as can be seen by evaluating on $x=e_{i},y=e_{j}$), hence SOS. Otherwise, we may scale $Q$ so that its diagonal coefficients become $1$. Then $Q$ becomes a monic symmetric biquadratic form, which by the first part is SOS. Scaling back preserves the SOS property. Therefore $Q$ is SOS. ∎
5.6 A Conjecture on $3\times 3$ Symmetric Forms
Based on these observations, we propose the following conjecture.
@@ 5.6 A Conjecture on $3\times 3$ Symmetric Forms
Theorem 5.3 provides a complete characterization of the PSD polytope for $3\times 3$ monic symmetric forms, and Theorem 5.5 establishes that every such form is SOS. The explicit SOS decompositions presented here, along with the known sos rank of $P_{0}$ (four) and the numerical evidence for $P_{4}$ (five), suggest that the sos rank may be bounded by five. Resolving this conjecture would provide a precise quantitative understanding of the SOS complexity for this class.
@@ 6 Further Remarks
A particularly promising direction for future research lies in the study of symmetric biquadratic forms. By focusing on forms that are symmetric in the variables $\mathbf{x}$ and/or $\mathbf{y}$, we navigate away from known pathological counterexamples like Choi’s, which lacks such symmetry. The special case of $3\times 3$ monic symmetric biquadratic forms, fully characterized in Theorem 5.2 and whose PSD polytope is described in Theorem 5.3, serves as an ideal testbed. Theorem 5.5 establishes that every psd $3\times 3$ symmetric biquadratic form is sos, resolving the existence question for this class. Conjecture 5.6, which posits that every such form has sos rank at most five, is a natural and significant next target. A positive resolution would provide a precise quantitative understanding of the SOS complexity for this class, marking substantial advancement on a modern analogue of Hilbert’s 17th problem for a structured class of polynomials.
Remark on extending Theorem 5.5 to $m,n>3$. While Theorem 5.5 establishes that every psd $3\times 3$ symmetric biquadratic form is SOS, extending this result to general $m\times n$ dimensions presents significant challenges. For $3\times 3$, the PSD cone of monic symmetric forms is a convex polyhedron defined by four linear inequalities (Theorem 5.3), allowing a vertex-by-vertex verification that all extreme points are SOS. For $m,n>3$, however, the PSD conditions in Theorem 5.2 include nonlinear constraints (e.g., $1-ab/c\geq 0$ when $c\neq 0$), and the PSD set is no longer a polytope. Moreover, the M-eigenvalue characterization becomes more intricate, and explicit SOS constructions at the boundary of the PSD cone are not readily generalizable. Even if every psd symmetric form were SOS for all $m,n$, the minimal number of squares (the SOS rank) may grow with $m$ and $n$. Resolving the extension to higher dimensions remains an open problem, requiring new techniques from convex algebraic geometry and tensor spectral theory.
Acknowledgment We are thankful to Professor Giorgio Ottaviani, who proposed to study symmetric biquadratic forms. This work was partially supported by Research Center for Intelligent Operations Research, The Hong Kong Polytechnic University (4-ZZT8), the National Natural Science Foundation of China (Nos. 12471282 and 12131004), the R&D project of Pazhou Lab (Huangpu) (Grant no. 2023K0603), the Fundamental Research Funds for the Central Universities (Grant No. YWF-22-T-204), and Jiangsu Provincial Scientific Research Center of Applied Mathematics (Grant No. BK20233002).

### v2

@@ 4.3 The PSD Cone is a Convex Polyhedron
Let $P$ be an $m\times n$ monic symmetric biquadratic form as in (6). Then the PSD set of parameters $(a,b,c)$ is a convex polyhedron (a tetrahedron) in $\mathbb{R}^{3}$ with vertices
$V_{1}$ $=(1,1,1),$
$V_{2}$ $=\left(-\frac{1}{m-1},\;-\frac{1}{n-1},\;\frac{1}{(m-1)(n-1)}\right),$
$V_{3}$ $=\left(-\frac{1}{m-1},\;1,\;-\frac{1}{m-1}\right),$
$V_{4}$ $=\left(1,\;-\frac{1}{n-1},\;-\frac{1}{n-1}\right).$
Consequently, condition (v) of Theorem 4.2 is redundant; the PSD set is exactly the convex hull of these four vertices.
@@
The inequalities (i)–(iv) of Theorem 4.2 are linear in $a,b,c$ and define a polyhedron in $\mathbb{R}^{3}$. Solving the systems of three equations obtained by setting any three of these inequalities to equality yields precisely the four vertices $V_{1},V_{2},V_{3},V_{4}$. One checks that each vertex satisfies all four inequalities, hence belongs to the polyhedron. Moreover, the polyhedron is bounded because the inequalities imply bounds on $a,b,c$ independently (for example, from (i) and (ii) together with (iii) one obtains $a,b,c\in[-1,1]$ when $m,n\geq 2$). Thus the polyhedron is a tetrahedron.
It remains to show that condition (v) is automatically satisfied for every point in this tetrahedron. Since the tetrahedron is convex, any point $(a,b,c)$ in it can be written as a convex combination of the vertices. In the next subsection we will exhibit explicit SOS decompositions for each vertex, proving that all vertices correspond to PSD (in fact SOS) forms. Because the set of PSD forms is a convex cone, any convex combination of the vertices also yields a PSD form. Hence, by Theorem 4.2, condition (v) must hold for every such point. Therefore condition (v) is redundant, and the PSD set coincides with the tetrahedron. ∎
4.4 SOS Decompositions of the Vertices
The four vertices admit simple explicit SOS representations.
@@ 4.4 SOS Decompositions of the Vertices
For any $m,n\geq 2$, the monic symmetric biquadratic forms corresponding to the vertices $V_{1},V_{2},V_{3},V_{4}$ are SOS. Explicitly,
$P_{V_{1}}({\bf x},{\bf y})$ $=\Bigl(\sum_{i=1}^{m}x_{i}\Bigr)^{2}\Bigl(\sum_{j=1}^{n}y_{j}\Bigr)^{2},$
$P_{V_{2}}({\bf x},{\bf y})$ $=\frac{1}{(m-1)(n-1)}\sum_{1\leq i<k\leq m}\sum_{1\leq j<l\leq n}\bigl((x_{i}-x_{k})(y_{j}-y_{l})\bigr)^{2},$
$P_{V_{3}}({\bf x},{\bf y})$ $=\frac{1}{m-1}\sum_{1\leq i<k\leq m}\Bigl((x_{i}-x_{k})\sum_{j=1}^{n}y_{j}\Bigr)^{2},$
$P_{V_{4}}({\bf x},{\bf y})$ $=\frac{1}{n-1}\sum_{1\leq j<l\leq n}\Bigl(\Bigl(\sum_{i=1}^{m}x_{i}\Bigr)(y_{j}-y_{l})\Bigr)^{2}.$
Proof.
Direct expansion shows that each expression equals the monic symmetric form with the respective parameters $(a,b,c)$. Each right‑hand side is a sum of squares of bilinear forms. ∎
4.5 Every PSD Symmetric Biquadratic Form is SOS
Combining the convexity of the PSD cone with the SOS property of the vertices yields the main result of this section.
@@ 4.5 Every PSD Symmetric Biquadratic Form is SOS
Every PSD monic symmetric biquadratic form (for any $m,n\geq 2$) is a sum of squares of bilinear forms. Consequently, every PSD symmetric biquadratic form is SOS.
@@
By Theorem 4.3, the set of coefficients $(a,b,c)$ giving a PSD monic symmetric form is the tetrahedron with vertices $V_{1},V_{2},V_{3},V_{4}$. Proposition 4.4 shows that each vertex corresponds to an SOS form. Since the mapping from $(a,b,c)$ to the form $P$ is linear, any convex combination of the vertices yields the same convex combination of the associated forms. The set of SOS forms is a convex cone; therefore any convex combination of SOS forms is again SOS. Hence every point in the tetrahedron corresponds to an SOS form, proving the first statement.
Now let $Q$ be an arbitrary $m\times n$ symmetric biquadratic form that is PSD. By symmetry, all its diagonal coefficients $q_{ijij}$ are equal. If they are all zero, then evaluating $Q$ on $x=e_{i},y=e_{j}$ shows that $Q$ is identically zero, which is trivially SOS. Otherwise, we may scale $Q$ so that its diagonal coefficients become $1$. The scaled form is then a monic symmetric biquadratic form, which by the first part is SOS. Scaling back preserves the SOS property. Thus $Q$ is SOS. ∎
4.6 The $3\times 3$ Case and a Conjecture on SOS Rank
For illustration we restate the $3\times 3$ case, which was historically the first nontrivial instance where the SOS property was unknown.
Corollary 4.6.
Every $3\times 3$ PSD symmetric biquadratic form is SOS.
Proof.
This is the special case $m=n=3$ of Theorem 4.5. ∎
The explicit SOS decompositions given in Proposition 4.4 for the vertices, together with the known SOS rank of the interior point $P_{0}({\bf x},{\bf y})=(x_{1}^{2}+x_{2}^{2}+x_{3}^{2})(y_{1}^{2}+y_{2}^{2}+y_{3}^{2})$ (which is four), suggest that the SOS rank of any $3\times 3$ symmetric PSD form might be bounded by five.
@@ 4.6 The $3\times 3$ Case and a Conjecture on SOS Rank
4.7 Illustrative Special Cases
For concreteness we record the explicit inequalities and vertices for two cases that originally motivated the general result.
4.7.1 The $4\times 3$ Case
Theorem 4.8.
Let $P$ be a $4\times 3$ monic symmetric biquadratic form as in (6) with $m=4$, $n=3$. Then $P$ is PSD if and only if the coefficients $(a,b,c)$ satisfy the four linear inequalities
$1-b+3(a-c)$ $\geq 0,$ (9)
$1-a+2(b-c)$ $\geq 0,$ (10)
$1-a-b+c$ $\geq 0,$ (11)
$1+3a+2b+6c$ $\geq 0.$ (12)
These inequalities define a tetrahedron with vertices
$V_{1}=(1,1,1),\quad V_{2}=\left(-\tfrac{1}{3},-\tfrac{1}{2},\tfrac{1}{6}\right),\quad V_{3}=\left(-\tfrac{1}{3},1,-\tfrac{1}{3}\right),\quad V_{4}=\left(1,-\tfrac{1}{2},-\tfrac{1}{2}\right).$
Proof.
Specialize Theorem 4.3 to $m=4$, $n=3$. ∎
4.7.2 The $4\times 4$ Case
Theorem 4.9.
Let $P$ be a $4\times 4$ monic symmetric biquadratic form as in (6) with $m=n=4$. Then $P$ is PSD if and only if the coefficients $(a,b,c)$ satisfy the four linear inequalities
$1-b+3(a-c)$ $\geq 0,$ (13)
$1-a+3(b-c)$ $\geq 0,$ (14)
$1-a-b+c$ $\geq 0,$ (15)
$1+3a+3b+9c$ $\geq 0.$ (16)
These inequalities define a tetrahedron with vertices
$V_{1}=(1,1,1),\quad V_{2}=\left(-\tfrac{1}{3},-\tfrac{1}{3},\tfrac{1}{9}\right),\quad V_{3}=\left(-\tfrac{1}{3},1,-\tfrac{1}{3}\right),\quad V_{4}=\left(1,-\tfrac{1}{3},-\tfrac{1}{3}\right).$
Proof.
Specialize Theorem 4.3 to $m=n=4$. ∎
The SOS decompositions for these vertices are exactly the ones given in Proposition 4.4 with the corresponding $m,n$. Consequently, every $4\times 3$ or $4\times 4$ PSD symmetric biquadratic form is SOS, as guaranteed by Theorem 4.5.
@@ 5 Further Remarks
A particularly promising direction for future research lies in the study of symmetric biquadratic forms. By focusing on forms that are symmetric in the variables $\mathbf{x}$ and/or $\mathbf{y}$, we navigate away from known pathological counterexamples like Choi’s, which lacks such symmetry. Theorem 4.5 establishes that every PSD symmetric biquadratic form (for any dimensions $m,n$) is SOS, completely resolving the existence question for this class. The PSD cone of monic symmetric forms is always a convex polyhedron (a tetrahedron) whose vertices admit explicit SOS decompositions. Conjecture 4.7, which posits that every $3\times 3$ symmetric PSD form has SOS rank at most five, is a natural and significant next target. A positive resolution would provide a precise quantitative understanding of the SOS complexity for this class, marking substantial advancement on a modern analogue of Hilbert’s 17th problem for a structured class of polynomials.
Acknowledgement We are thankful to Professor Giorgio Ottaviani, who proposed to study symmetric biquadratic forms. This work was partially supported by Research Center for Intelligent Operations Research, The Hong Kong Polytechnic University (4-ZZT8), the National Natural Science Foundation of China (Nos. 12471282 and 12131004), the R&D project of Pazhou Lab (Huangpu) (Grant no. 2023K0603), the Fundamental Research Funds for the Central Universities (Grant No. YWF-22-T-204), and Jiangsu Provincial Scientific Research Center of Applied Mathematics (Grant No. BK20233002).